# Image Generation from Contextually-Contradictory Prompts

### v1

@@ 3.2. Prompt Decomposition
Figure 4. The effect of the object removal strategy. Each image is generated from a prompt that initially describes only the background (“a white sandy beach,” “a desert”). At the timestep indicated below each column, the foreground object (“a monkey,” “a snowman”) is added to the prompt. Removing the object from the initial prompt, without using a placeholder, significantly affects its apparent scale. Early insertion allows the model to allocate more spatial area, while late insertion confines the object to limited regions within the established layout.
3.3. Stage-aware Prompt Injection
Given a sequence of proxy prompts ${p_{1},p_{2},...,p_{n}}$, their corresponding timestep intervals ${I_{1},I_{2},...,I_{n}}$, and a text-to-image (T2I) diffusion model, we condition the model using different prompts throughout the denoising process. At each timestep $t$, we apply the prompt $p_{i}$ such that $t\in I_{i}$. By aligning each proxy prompt with a specific timestep interval, the denoising process is guided at each stage by concepts relevant to the level of detail emerging at that point. This enables the gradual construction of the image while avoiding conflicts with the model’s learned priors.
This injection mechanism integrates seamlessly into existing inference pipelines without any architectural modifications, and is compatible with a range of pretrained diffusion models. Differences in denoising schedules across models are handled entirely during prompt decomposition (Section 3.2). We show examples with SD3 in the appendix (see Figure 11).
“A camping tent is inside a bed room”
@@ 3.3. Stage-aware Prompt Injection
too early too late Ours
Figure 5. Impact of proxy object timing on generated image attributes. Changing the target object too early fails to disentangle the desired attribute, while modifying it too late prevents effective binding. Attribute control is most successful when applied within the correct time interval. In the top row, the initial proxy prompt is “A pillow fort in a bedroom”, while in the bottom row, the proxy prompt is “A mother duck guards three ducklings”.
4. Experiments and Results
In this section, we conduct qualitative (Section 4.1) and quantitative (Section 4.2) comparisons to assess the effectiveness of our method, along with ablation study (Section 4.3) to evaluate the contribution of individual components. We also present some limitations of our method (Section 4.4).
Implementation Details.
We use the state-of-the-art text-to-image model FLUX.1 [dev] (Labs, 2024) and GPT-4o as LLM to demonstrate the effectiveness of our approach. Inference is performed using $T=50$ sampling steps. In all experiments, we guide the LLM to limit the number of proxy prompts generated during decomposition to a maximum of three per input prompt. We use the baselines hyperparameters as appeared in their original papers or implementation. To ensure a fair comparison, we also use $T=50$ sampling steps for all baseline methods.
Baselines
Since our method builds upon FLUX.1 [dev], we directly compare our results against this base model to clearly demonstrate the improvements achieved by our approach. Additionally, we include a comparison to the training-free method R2F (Park et al., 2024), which similarly employs a large language model (LLM) to alternate between different prompts during generation. As R2F is implemented based on SD3 (Esser et al., 2024), we also compare our results to this model. Lastly, we compare against the fine-tuned model Ella (Hu et al., 2024), which is currently publicly available with SD 1.5 as its base model.
Datasets
We evaluate our method using three datasets: Whoops! (Bitton-Guetta et al., 2023), Whoops-Hard, and ContraBench. Whoops! consists of 500 prompts paired with commonsense-defying images, designed to test visual reasoning and compositional understanding. While relevant to our task, many of its prompts are relatively easy for modern T2I models and fail to consistently reveal their limitations. To address this, we curate a more challenging dataset Whoops-Hard, a subset comprising 100 particularly difficult prompts from Whoops! (Bitton-Guetta et al., 2023). This challenging benchmark provides a stronger basis for evaluating text-to-image generation given the current performance of state-of-the-art methods. Furthermore, to test more challenging cases of semantic alignment, we introduce a new benchmark ContraBench, a curated set of 40 prompts crafted to induce strong contextual contradictions, including background-object mismatches, semantic tension, and scale conflicts. ContraBench serves as a focused, high-difficulty benchmark for evaluating text-image alignment on prompts that contain contextual contradictions. We will make both ContraBench and the Whoops-Hard subset publicly available.
@@ 4. Experiments and Results
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ella SD3 R2F FLUX Ours
Figure 6. Our method effectively resolves contextual contradictions across diverse forms, including inconsistent backgrounds, unnatural color assignments, and implausible actions. Compared to prior approaches, our results exhibit higher semantic coherence and more accurate prompt grounding under contradictory prompts.
4.1. Qualitative Results
Comparisons with baselines.
@@ 4.1. Qualitative Results
Figures 6 and 9 present qualitative comparisons on the Whoops! and ContraBench datasets. Across both benchmarks, the baseline methods FLUX, SD3, R2F, and Ella consistently exhibit characteristic failure modes when handling prompts with contextually contradictory concepts. For instance, as shown in Figure 6, our method is the only one to successfully generate scenarios such as a blue Shrek or a monkey juggling tiny elephants.
@@
For SD3 and FLUX, these prompts conflict with their learned priors, leading to failures in generating text-aligned images. Ella was originally designed for long and complex prompts, not specifically for those containing contextual contradictions, and therefore struggles with challenging contradiction cases. The training-free baseline R2F alternates between prompts at predefined timesteps, a strategy originally designed for attribute binding rather than addressing contextual contradictions. While it can reinforce rare concepts, this approach does not align prompt content with the stages at which semantic features emerge during denoising. As a result, it often produces hybrid concepts that merge incompatible elements from conflicting objects (see the owl and SpongeBob in Figure 9).
@@
In contrast, our method produces semantically coherent outputs by introducing proxy prompts at denoising stages where the corresponding visual features are expected to emerge. This alignment enables the model to effectively handle conflicting concepts. Across both Whoops! and ContraBench, our approach consistently generates contextually contradictory elements with high accuracy.
To demonstrate the robustness of our approach, we provide a comparison across several random seeds in Figure 10. The same seeds are used for all methods, showing that our method consistently outperforms the baselines.
Improve realism
@@
Figure 7 demonstrates how our method generates photorealistic and semantically coherent images for prompts with atypical attribute combinations. In contrast, FLUX often defaults to cartoon-like renderings, likely due to priors associating fantastical elements with imaginative visual styles. By using contradiction-free proxy prompts, our method enables the model to produce realistic outputs without explicitly requiring photorealism in the prompt.
@@
Figure 7. FLUX tends to generate realistic images by default. However, when given unrealistic prompts, it often produces cartoon-like samples. In contrast, our method, which gradually resolves such prompts through coherent proxy stages, consistently generates realistic and semantically aligned images.
4.2. Quantitative Results
Our method focuses on solving alignment with challenging and contradictory textual prompts. Therefore, we evaluate its performance by measuring prompt alignment. To do so, we use the GPT-4o vision-language model (VLM). The model is prompted to assign a score from 1 to 5 for each generated image, based on how well the image aligns with the given textual prompt. The exact evaluation prompt is provided in the appendix (See Table 6). For each prompt, we generate outputs using three fixed random seeds, which are applied consistently across all evaluated methods. The average score across the three seeds is then scaled to a range between 20 and 100 by multiplying by 20.
As can be seen in Table 1, our method outperforms the baselines across all three benchmarks. Notably, using SD3 as the base model yields better prompt alignment than the state-of-the-art FLUX model under contradictory prompt settings, while FLUX demonstrates superior visual quality (See Figure 9). In contrast, our method, built on top of FLUX, achieves the highest prompt adherence while preserving its strong visual quality. This results in the best overall performance, combining high-fidelity generation with robust alignment to challenging prompts.
Table 1. Quantitative evaluation on the Whoops dataset using the GPT-4o vision-language model. We report average alignment, where alignment reflects how well the image matches the prompt semantics, without looking at the quality of the image
Benchmarks
Models Whoops Whoops-Hard ContraBench
Ella 69.09 45.19 55.16
SD3.0 82.63 55.73 57.5
R2F (SD3.0) 83.50 57.06 59.16
FLUX 78.85 44.3 57.16
Ours (FLUX) 85.10 62.13 66.16
Table 2. User study results. Win rates of our method in text-image alignment and image quality, compared against each baseline method.
Ella SD3.0 R2F (SD3.0) FLUX
Alignment $81\%$ $70\%$ $75\%$ $81\%$
Quality $74\%$ $72\%$ $68\%$ $63\%$
User study
VLM-based automatic metrics, such as those shown in Table 1, often fail to detect subtle semantic inconsistencies, particularly when evaluating generations conditioned on complex or contradictory prompts. Moreover, these models do not adequately assess the visual quality of the generated images. To address these limitations, we conducted a user study aimed at evaluating both prompt adherence and visual quality from a human perspective. We randomly sampled 24 prompts from the Whoops! and ContraBench benchmarks. For each prompt, participants were shown two images: one generated by our method and the other by one of the baselines. They were asked to answer two questions: (1) which image most accurately reflected the prompt, and (2) which image had better visual quality.
In total, we collected responses from 61 users, resulting in 1,464 individual evaluations. The results of the user study are summarized in Table 2, which reports the win rate of our method when compared pairwise against each baseline. These results highlight the superiority of our approach in handling contextually contradictory prompts, achieving both better prompt alignment and visual quality compared to baseline approaches.
Refer to caption Refer to caption Refer to caption
Figure 8. Failure cases from Whoops!.
4.3. Ablation Studies
To evaluate the contribution of each component in our method, we perform an ablation study on the Whoops-Hard benchmark by selectively removing or modifying individual elements of the method, and present the results in Table 3. Each configuration isolates a specific design choice to quantify its effect on alignment performance. We observe that using in-context examples significantly improves the LLM’s ability to decompose and interpret contextually contradictory prompts, resulting in better text-image alignment. Removing the explanation requirement from the LLM impairs its reasoning capabilities and leads to a notable performance drop, suggesting that requiring the LLM to generate explicit explanations helps it make more coherent semantic decisions, ultimately improving prompt decomposition. Additionally, using only two proxy prompts, underperforms compared to the full method, highlighting the importance of introducing prompt information at targeted stages for robust alignment.
4.4. Limitations
In Figure 8, we present three failure cases of our method on examples from the Whoops! benchmark. Since our approach relies on guiding the model through text alone, it cannot control properties that the underlying model inherently struggles with—such as generating specific quantities or enforcing precise orientations.
Table 3. Ablation study on the Whoops-Hard Benchmark. We evaluate the impact of key components of our prompt decomposition method by: (1) removing in-context examples, (2) removing the explanation requirement, and (3) restricting the decomposition to at most two proxy prompts. Results show a significant drop in alignment when any of these components are removed.
static w/o in-context w/o reasoning 2 proxy Full
Alignment 44.3 48.0 46.46 50.73 62.13
5. Conclusions
We have introduced a training-free framework for resolving contextual contradictions in text-to-image generation – that is, cases where seemingly plausible prompts are internally inconsistent because of strong, hidden biases in the model. Our method identifies and strategically separates such contradictions over time, enabling faithful rendering of prompts that would otherwise yield semantically inconsistent outputs.
At its core, our approach leverages the broad real-world understanding embedded in large language models to reason about the semantic plausibility of prompts. Rather than modifying the training data or fine-tuning the model, we harness this reasoning capacity at inference time to decompose prompts into a sequence of contextually coherent stages. These proxy prompts guide the generative process in alignment with the model’s prior expectations, helping it circumvent conflicts and maintain semantic fidelity.
Extensive experiments demonstrate that our stage-aware decomposition substantially improves alignment and image quality – especially for prompts that challenge the model’s default associations. Future directions include refining our method using VLM-generated responses to improve proxy prompts generation dynamically during the generation process, and applying our method to text-to-video generation models.
Appendix A Appendix

### v2

@@ 5 Experiments and Results
Figure 6: Early insertion of the foreground allows the model to allocate more space to it, while late insertion confines the object within the existing layout, making it appear smaller (e.g., the snowman introduced at Step 5). Step labels indicate when the foreground object is introduced.
Implementation Details.
We use FLUX.1 [dev] [17] as the base T2I model and GPT-4o [1] for prompt decomposition. In all experiments, inference is performed using $T=50$ steps and the LLM is restricted to at most three proxies per prompt. Baseline hyperparameters follow their original papers or implementations, with $T=50$ steps for fair comparison. To further demonstrate robustness, we also report results with SD3.0 using the same LLM-generated proxy prompts and intervals.
Baselines.
We compare SAP against the following approaches: (1) base models FLUX-dev (denoted by FLUX) [17] and SD3.0 [11]; (2) R2F [22], a training-free method reported under three settings: SD3.0 (original), FLUX-schnell (official), and our reimplementation on FLUX; (3) Annealing Guidance [34], which trains a small MLP to predict the classifier-free guidance scale at each step; and (4) Ella [16], a fine-tuned model on SD1.5. (5) VL-DNP [8], a VLM-guided negative prompting method, evaluated using SD3 with Qwen2.5-VL-7B.
Datasets
We evaluate SAP using three datasets: Whoops! [7], Whoops-Hard, and ContraBench. Whoops! consists of 500 prompts paired with commonsense-defying images, designed to test visual reasoning and compositional understanding. While relevant to our task, many of its prompts are relatively easy for modern T2I models and do not consistently expose model limitations.
To address this, we curate Whoops-Hard, a subset of 100 particularly difficult prompts from Whoops!, providing a stronger benchmark for evaluating current state-of-the-art models. To further probe semantic alignment under contradictory conditions, we introduce ContraBench, a curated set of 40 prompts capturing contextual contradictions. The dataset was constructed in two steps: (1) ChatGPT generated candidate prompts based on the definition of contextual contradictions, and (2) human annotators filtered them to retain only those that clearly expressed contradictions. The full prompt lists for both datasets are provided in the Appendix.
5.1 Qualitative Results
“A camping tent is inside a bedroom.”
@@ 5.1 Qualitative Results
Step 1 Step 14 Ours
Figure 7: Effect of interval assignment. Introducing the full prompt too early fails to disentangle contextual contradictions, while introducing it too late alters only fine details. Top proxy: “A pillow fort in a bedroom”; bottom proxy “A mother duck guards three ducklings”.
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
@@
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Ella Annealing SD3 R2F FLUX R2F (schnell) R2F (FLUX) SAP
Figure 8: Qualitative comparison of SAP with baseline methods. Our method resolves contextual contradictions, whereas baselines struggle to produce text-aligned images. Additional examples are provided in the Appendix.
@@
Figure 8 presents qualitative comparisons on the Whoops! and ContraBench datasets (see Appendix for more results). Across both benchmarks, baseline methods consistently exhibit characteristic failure modes when handling contradictory prompts. In contrast, SAP successfully generates challenging cases such as a blue Shrek or a monkey juggling tiny elephants (Figure 8). In addition to FLUX, SAP also improves SD3 generations under contradictory prompts (Figure 9).
@@
For SD3 and FLUX, contradictory prompts expose conflicts with learned priors, resulting in prompt misalignment. Ella and Annealing Guidance, not designed for contradictions, perform less effectively on such cases. R2F alternates between prompts at predefined timesteps, a strategy designed for attribute binding rather than addressing contextual contradictions. While it can reinforce rare concepts, it does not align prompts with the stages at where semantic features emerge during denoising. As a result, it often produces hybrid concepts that merge incompatible elements from conflicting concept (see the bodybuilder in Figure 9, and the owl and SpongeBob results in the Appendix).
@@
In contrast, SAP produces semantically coherent outputs by introducing proxy prompts at denoising stages where corresponding features emerge. This enables effective handling of conflicting concepts. Across both benchmarks, SAP consistently generates coherent images for contradictory prompts. Robustness is further demonstrated in the Appendix, comparing generations across multiple seeds.
5.2 Quantitative Results
We evaluate prompt alignment using GPT-4o vision–language model (VLM). For each generated image, GPT-4o assigns a score from 1–5 based on adherence to the prompt. Scores are averaged across three fixed random seeds per prompt and scaled to 20–100. The evaluation prompt is provided in the Appendix.
As shown in Table 1, SAP outperforms all baselines across the three benchmarks. Between the base models, SD3.0 tends to yield stronger alignment under contradictions, while FLUX offers higher visual quality (Figure 8). SAP improves both backbones, enhancing prompt adherence while maintaining the visual fidelity of the underlying models (Figures 8 and 9 ).
SD3 $\text{R2F}_{\text{SD3}}$ $\textbf{SAP}_{\text{SD3}}$ Refer to caption Refer to caption Refer to caption “A grown man has a baby’s pacifier in his mouth.” Refer to caption Refer to caption Refer to caption “A bodybuilder balancing on pointe shoes.”
Figure 9: SAP is robust to the base model, as shown by the results obtained with $\text{SAP}_{\text{SD3}}$.
User study.
VLM-based metrics often miss subtle semantic inconsistencies and do not adequately assess image quality. To complement them, we conducted a user study evaluating both prompt adherence and overall visual appearance. We randomly sampled 24 prompts from the Whoops! and ContraBench benchmarks. For each prompt, participants compared two images, one generated by SAP and the other by a baseline, and answered: (1) which most accurately reflected the prompt, and (2) which had higher visual quality. In total, we collected responses from 61 users, yielding 1,464 individual evaluations. Table 2 summarizes the win rates of SAP against each baseline.
These results highlight the superiority of our approach in handling contextually contradictory prompts, achieving both stronger prompt alignment and higher visual quality.
5.3 Ablation Studies
We evaluate SAP through ablations that assess design choices in prompt decomposition and robustness under different conditions, such as timestep perturbations. Additional results on alternative LLMs and non-contradictory prompts are in the Appendix.
Table 1: Quantitative evaluation on various benchmarks using the GPT-4o vision-language model. We report average alignment, where alignment reflects how well the image matches the prompt semantics, independent of visual quality. SAP achieves the best results, regardless of the base model.
Benchmarks
Models Whoops Whoops- Contra-
Hard Bench
SD3.0 82.63 55.73 57.5
FLUX 78.85 44.3 57.16
Ella 69.09 45.19 55.16
Annealing 79.59 59.06 58.33
VL-DNP 82.79 56.26 58.83
R2F 83.50 57.06 59.16
$\text{R2F}_{\text{schnell}}$ 79.58 54.80 59.33
$\text{R2F}_{\text{FLUX}}$ 48.68 32.80 25.33
$\text{SAP}_{\text{SD3.0}}$ 85.87 64.40 65.33
SAP 85.10 62.13 66.16
Table 2: User study results. Win rates of SAP in text-image alignment and image quality, compared against each baseline method.
SD3 FLUX Ella Annealing R2F
Alignment $70\%$ $\% 8181\%$ $73\%$ $75\%$
Quality $72\%$ $63\%$ $74\%$ $79\%$ $68\%$
Table 3: Ablation on Whoops-Hard. We evaluate our prompt decomposition method by (1) removing in-context examples, (2) removing the explanation requirement, and (3) limiting decomposition to two proxy prompts.
static w/o in-context w/o reasoning 2 proxy Full
Alignment 44.3 48.0 46.46 60.26 62.13
Prompt decomposition components.
We conduct ablations on the Whoops-Hard benchmark, where each variant isolates a design choice to quantify its effect on alignment within our method (Table 3). In-context examples significantly improve the LLM’s ability to decompose contradictory prompts, leading to better text–image alignment. Removing the explanation requirement impairs reasoning and causes a notable drop, showing that generating explicit explanations encourages more coherent semantic decisions. Restricting decomposition to two proxies performs close to the full method, while allowing up to three proxies provides extra flexibility for harder cases and yields further gains.
Robustness to LLM-predicted timestep intervals.
Our method relies on LLM-predicted intervals to schedule proxy prompts, but these boundaries do not require exact placement. The earlier results (Figure 7) highlight that placing proxy prompts at the wrong stage of denoising (e.g., too early or too late) can harm alignment. Here we show that within the correct coarse stage, the method is robust to moderate boundary shifts. Specifically, we perturb interval boundaries while keeping the proxy prompts fixed, uniformly shifting them within windows of varying size. As shown in Table 4, small shifts of up to $\pm 1$ step (window=3) have almost no effect on alignment, and even larger shifts of up to $\pm 2$ steps (window=5) cause only minor degradation, despite representing a substantial perturbation relative to the full method effective range ($\sim$12 steps; see Appendix ). These results confirm that SAP is sensitive to the stage at which information is introduced, but largely insensitive to exact step boundaries within that stage.
Refer to caption Refer to caption Refer to caption
Figure 10: Failure cases of SAP.
Limitations.
In Figure 10, we present representative failure cases of our method on examples from the Whoops! benchmark. Since our approach relies on guiding the model through text alone, it cannot control properties that the underlying model inherently struggles with, such as generating specific quantities or enforcing precise orientations; a more detailed categorization and analysis of failure modes is provided in the supplementary.
Table 4: Effect of perturbing LLM-predicted timestep intervals. Boundaries are uniformly shifted within window $w$. $SAP_{w=i}$ denotes evaluation with window ii.
Benchmarks
Models Whoops Whoops- Contra-
Hard Bench
FLUX 78.85 44.3 57.16
SAP 85.10 62.13 66.16
$\text{SAP}_{w=3}$ 84.18 62.06 65.5
$\text{SAP}_{w=5}$ 81.46 58.46 62.5
6 Conclusions
We introduced a training-free framework for resolving contextual contradictions in text-to-image generation, cases where seemingly plausible prompts fail due to strong, hidden model biases. At its core, our method leverages the coarse-to-fine generation process to separate contradictions across denoising stages, enabling faithful rendering of prompts that would otherwise yield semantically inconsistent outputs. The introduction of proxy prompts steers the generative process in line with the model’s priors, enabling it to resolve conflicts and preserve semantic fidelity without the need for retraining or fine-tuning.
We argue that since our approach already leverages the broad world knowledge of vision–language models, integrating them more tightly with generative models holds promise for addressing contextual contradictions directly. As a next step, we plan to explore emerging compound architectures that combine VLMs and generative models, with the aim of understanding how to effectively harness them to resolve conflicts in open-ended generation.
Acknowledgments
This research was supported in part by the Israel Science Foundation (grants no. 2492/20 and 1473/24), Len Blavatnik and the Blavatnik family foundation.
Supplementary Material
7 Ethics Statement
Our work contributes to improving the semantic alignment of text-to-image models under contradictory or biased prompts. As a consequence, our method enhances users’ ability to control generative models and faithfully render contradictory concepts. While this provides positive benefits, such as reducing unintended biases and enabling more inclusive image generation, it also increases the potential for misuse, including the creation of harmful, misleading, or inappropriate content. As with any advance in generative modeling, these dual-use concerns highlight the importance of responsible deployment, safeguards, and continued ethical oversight to ensure that such improvements contribute positively to society.
8 Additional Results
Robustness across LLMs.
Since our framework hinges on LLM-driven prompt decomposition, we further examined its robustness under different language models. We evaluated both a proprietary model (GPT-4o) and a comparatively lightweight open-source alternative (LLaMA-3.1-8B-Instruct). While GPT-4o delivers the strongest performance, the smaller LLaMA-3.1-8B-Instruct still yields consistent improvements over the baseline (see Table 5).
Improved realism.
@@ 8 Additional Results
SAP generates photorealistic and semantically coherent images for prompts with atypical attribute combinations (Figure 11). In contrast, FLUX often defaults to cartoon-like renderings, even when photorealism is explicitly requested, revealing a contextual contradiction between fantastical content and realistic style. By using contradiction-free proxy prompts, SAP avoids these biases and produces realistic outputs regardless of whether photorealism is explicitly required in the prompt.
Non-contradictory prompts.
To ensure applicability in general text-to-image scenarios, we verify that our method does not negatively affect prompts without contextual contradictions. We find that including even a single non-contradictory in-context example is sufficient for the LLM to default to using the full prompt in such cases. We evaluate this behavior using GPT-4o alignment scores on the PartiPrompts-Simple benchmark, which contains simple, non-contradictory prompts (Table 6).
Additional qualitative comparisons.
Figures 13 and 14 present additional qualitative comparisons of our method, while Figure 15 shows results across multiple seeds.
Table 5: Performance of SAP when combined with different LLMs, comparing GPT-4o and Llama-3.1-8B-Instruct.
Benchmarks
Models Whoops Whoops- Contra-
Hard Bench
FLUX 78.85 44.3 57.16
$\text{SAP}_{\text{GPT4o}}$ 85.10 62.13 66.16
$\text{SAP}_{\text{Llama3.1}}$ 80.52 59.53 61.16
Table 6: Alignment performance on the PartiPrompts-simple benchmark, which contains simple, non-contradictory prompts. Scores are computed using GPT-4o vision-language model. Our method achieves comparable performance to the base model, indicating no degradation on regular prompts.
Models PartiPrompts-simple
FLUX 93.46
SAP 93.06
@@
Figure 11: FLUX tends to generate realistic images by default. However, when given unrealistic prompts, it often produces cartoon-like samples. In contrast, our method, which gradually resolves such prompts through coherent proxy stages, consistently generates realistic and semantically aligned images.
Failure Mode Analysis.
We analyze failure cases of SAP and group them into three categories, summarized in Table 7. The majority of failures arise from incorrect proxy prompt assignment, where the selected proxy does not adequately preserve the intended structure or semantics of the target concept. In many such cases, manually correcting the proxy resolves the contradiction, as illustrated in Figure 12 (left), indicating that these failures stem from limitations of the LLM rather than the proposed framework, and can potentially be mitigated. A second group of failures stems from inherent limitations of the underlying diffusion model, such as difficulty generating an empty pool or rendering unusual symbolic structures (e.g., compass directions), which persist even with correct prompt decomposition (Figure 12, right). Finally, only a small fraction of failures is attributed to suboptimal interval assignment, consistent with our findings that SAP is robust to moderate variations in timestep scheduling.
Table 7: Breakdown of SAP failure cases by root cause.
Interval Proxy Model
assignment assignment limitation
Failure cases 3.3% 63.3% 33.3%
Bad proxy / Corrected Model limitation
Refer to caption Refer to caption Refer to caption Refer to caption
“A man is welding
without a mask”
“A man jumping
into an empty
swimming pool.”
“A compass with
North South South
West points”
Figure 12: Failure cases of SAP. Left: a proxy assignment failure, where an unsuitable proxy leads to incorrect generation; manually correcting the proxy (right image in pair) resolves the contradiction. Right: failures due to inherent limitations of the underlying diffusion model, such as difficulty generating an empty pool or rendering unusual symbolic structures (e.g., a compass with incorrect directions).
Computational Overhead.
SAP introduces modest overhead due to a single LLM inference for prompt decomposition, performed once per prompt prior to diffusion, and a lightweight additional embedding pass. As this step occurs only once and is independent of the denoising process, it does not significantly impact overall generation time.
9 LLM Instruction for Prompt Decomposition
Tables 8 and 9 detail the full LLM instruction used for our method’s decomposition, along with the corresponding in-context examples. In a single inference pass, our method detects contextual contradictions, generates proxy prompts, and assigns timestep intervals.
10 Provided Benchmarks
We describe the construction of ContraBench and Whoops-Hard in the main text. Here, we provide the full lists of prompts for these benchmarks in Table 11 and Table 12, respectively.
11 VLM Evaluation
We utilize GPT-4o to assess alignment between prompts and their generated images. The instruction prompt provided to the VLM is shown in Table 13.
Ella Annealing SD3 R2F Flux SAP
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A balloon is lifting up a package.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A coocoo clock with a horse popping out.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“An owl is perched upside down on a branch.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A photorealistic photo of SpongeBob SquarePants dancing ballet.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“Children in Africa are building a snowman.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A cruise ship parked in a bathtub.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A fortune teller predicts the future with a basketball.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A Viking warrior flying an airplane.”
Figure 13: Qualitative comparison. Our method consistently generates text-aligned images for contextually contradicting prompts.
Ella Annealing SD3 R2F Flux SAP
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“An Egyptian pyramid is constructed using modern construction gear.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A cat balancing a skyscraper on its nose.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A cowboy swimming competitively in an Olympic pool.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A knight in chess is a unicorn.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A man giving a piggyback ride to an elephant.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A butterfly is in a bee’s hive.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“An eagle is swimming under-water.”
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A photorealistic image of a bear ironing clothes in a laundry room”
Figure 14: Qualitative comparison. Our method consistently generates text-aligned images for contextually contradicting prompts.
Ella Annealing SD3 R2F Flux SAP
Seed 1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Seed 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Seed 3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“a professional boxer does a split.”
Seed 1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Seed 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Seed 3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A woman writing with a dart.”
Seed 1 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Seed 2 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Seed 3 Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
“A man riding a jet ski through the desert.”
Figure 15: Qualitative comparison across multiple seeds. Our method consistently generates text-aligned images for contextually contradicting prompts.